# Concept-1K: A Novel Benchmark for Instance Incremental Learning

### v2

@@ Appendix A Evaluation Metrics
Table 6: The statistics of the PLMs used in this paper. $\dagger$: Non-embedding parameters according to Biderman et al. (2023).
@@
Appendix B Experimental Settings
@@
Backbones. We use the Pythia suite Biderman et al. (2023) and the popular open-source models including LLaMa and Vicuna for experiments. Pythia is based on GPT-NeoX Black et al. (2022), which contains 8 model sizes and 154 pre-training checkpoints, enabling research in interpretability and learning dynamics. The statistics of the 9 PLMs used in this paper are summarised in Table 6. We download the pre-trained weights from Huggingface (Wolf et al., 2019).
Datasets. We split the 1023 concepts in Concept-1K into 10 groups for IIL. Specifically, they sorted the concepts alphabetically and shuffled the order with random seed 1. The first group contains 105 concepts, while the others contain 102 concepts.
Baseline Methods. The introduction of the baseline methods is as follows:
  • SEQ: SEQ refers to sequential fine-tuning. It can be considered as the lower bound of incremental learning.
  • REPLAY: Experience replay stores representative old samples and jointly optimizes old and new samples when learning new tasks. It is a practical and popular technique in incremental learning.
  • EWC Kirkpatrick et al. (2017): EWC is a regularization-based method. The weight of each parameter is determined by the diagonal of the Fisher information matrix. The weight of the regularization loss is searched in {5×$\times 10^{3}$, 1×$\times 10^{4}$, 5×$\times 10^{4}$, 1×$\times 10^{5}$, 1×$\times 10^{6}$, 1×$\times 10^{7}$}.
  • LAMOL Sun et al. (2020): LAMOL trains PLMs with question-answering and generative objectives. It generates pseudo-samples before learning each new task for data replay. The weight of the generation loss is $\lambda=0.25$. The proportion of pseudo samples is $\gamma=0.20$. There are two variants of LAMOL, including LAMOL_t and LAMOL_g. The difference is whether a task-specific token is used for generation.
  • L2KD Chuang et al. (2020): L2KD adds a knowledge distillation target based on LAMOL. The teacher model is trained from scratch. We implemented the word-level variant since it performs best on text classification tasks.
  • LAMOL_KD Zheng et al. (2023): LAMOL_KD utilizes knowledge distillation based on LAMOL_t. Unlike L2KD, the teacher model in LAMOL_KD is trained on all previous tasks. The new data are used to learn the LAMOL objectives, and pseudo data are used for word-level knowledge distillation as a regularization term.
  • PCLL Zhao et al. (2022): PCLL combines the idea of variational autoencoders and word-level knowledge distillation with the objective of LAMOL.
  • LFPT5 Qin and Joty (2022): LFPT5 only learns soft prompts for each new task. The training objective is the same as LAMOL. The number of soft prompt tokens is 10.
  • LoRA Hu et al. (2021): LoRA only trains a small proportion of parameters of PLMs. We set the rank $r=8$ and the scaling parameter $\alpha=16$. We use the implementation of LoRA from the PEFT library Mangrulkar et al. (2022).
We note that some IL methods are not compared since they are not applicable in the IIL scenario. For example, Progressive Prompt Razdaibiedina et al. (2023) requires task ID during both training and inference stages. VAG Shao et al. (2023) requires storing the vocabulary of class labels and do not apply to the generation tasks without class labels. Besides, prompt-based IL methods such as L2P Wang et al. (2022) also do not apply to generation tasks.
@@
Implementation Details. The maximum input and output lengths are set as 32 and 10, respectively. The batch size is 32. The learning rate of PLMs is $1\times 10^{-5}$. We use AdamW Loshchilov and Hutter (2018) optimizer. We use A800 GPUs for our experiments. Each experiment is repeated three times, and the average and standard deviations are reported. We search for the best hyper-parameters for each baseline method.
Figure 3: The accuracy of different concepts in the first task after sequential finetuning Pythia-2.8B on the 10 tasks of Concept-1K.
Appendix C Additional Experimental Results and Analysis
C.1 The Role of Model Scale, Pretraining, and Buffer Size
We provide the detailed result of Figure 2 in Table 7, 8, 9, 10, 11, 12, 13, and 14.
C.2 The performance of SOTA methods
The backbone model is Pythia-410M. We provide the step-wise average accuracy on training and test set in Figure 4 (a) and (b). The results demonstrate a large gap between SOTA IL methods and data replay. We also provided a comparison between SOTA methods in terms of memorization accuracy, generalization accuracy, memorization forgetting, and generalization forgetting in Figure 4 (c), (d), (e), and (f). In addition, from Figure 4 (g), we find that the training loss of the prompt-tuning-based method LFPT5 can not decrease to a small value. It indicates that only finetuning prompts are not practical for learning new knowledge, which is also consistent with the result of LoRA.
C.3 In-Context Learning
Apart from the analysis in the main paper, Table 5 shows that the smallest PLM (Pythia-70M) achieve high accuracy under the “1 shot” and “Same Know.” settings. The reason is that small PLMs simply copy the output in the demonstration instance as the final output. Under the “5 shot” and “Same Know.” settings, the accuracy of Pythia-70M is only 15.02%.
We use the model “gpt-3.5-turbo” and “gpt-4-turbo” for GPT3.5 and GPT4, respectively. Due to the high cost of evaluating on the whole concept-1K, we randomly sample 500 instances for GPT3.5 and GPT4. The output and the target of GPT3.5 and GPT4 on the 500 instances are provided in the supplementary material. Table 5 shows that GPT4 achieves 86.20% under the “5 shot” and “Same Know.” settings. It indicates that the training and the test instance of Concept-1K share the same knowledge.
C.4 Concept-Level Accuracy on Concept-1K
We analyze the memorization accuracy of the concepts in the first task after IIL. The backbone model is Pythia-2.8B, and it is sequential finetuning on Concept-1K. From Figure 3, we have the following observation:
  • The concepts such as “Cultural Competence”, “Augmented Reality Displays”, “Geopolitical Risks” that achieve high accuracy are broad, interdisciplinary topics that are likely rich in diverse data sources and content. These subjects span various fields and are likely to have clear definitions, significant public interest, and extensive coverage in the pretraining datasets, facilitating high memorization accuracy.
  • The lowest accuracy is observed in highly technical, cutting-edge, or niche fields such as “Spatial Audio”, “Biomaterials”, “Sustainable Tech”, and “Space Tourism”. These topics are likely challenging due to rapidly evolving information, specialized knowledge requirements, and potentially limited representation in available datasets.
In summary, the tendency from high to low accuracy reflects the transition from broad, interdisciplinary topics with abundant and diverse data to more specialized, technical, or emerging fields with potentially limited data exposure. This suggests that the PLMs’ memory accuracy is influenced by the richness and diversity of information available in its training data and the clarity and stability of the concepts themselves. Additionally, despite their significance, emerging technologies and areas of innovation might be harder for PLMs to learn accurately due to the rapid evolution of knowledge and possibly less structured data.
@@ C.4 Concept-Level Accuracy on Concept-1K
Table 16: (Continual) Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance.
Domain Concept Triplet Training and Test Input Target Output
What kind of process is Groundwater Recharge?
What are essential for the process of Groundwater Recharge?
What does Groundwater Recharge lead to regarding water tables?
Why is groundwater recharge important?
What is the primary factor leading to Sea Level Rise?
What group of professionals study Sea Level Rise?
What is a significant impact of Sea Level Rise on natural habitats?
What instrument is used to measure sea level rise?
What strategy addresses Sea Level Rise?

### v3

@@ E.3 Trend criterion
Table 10: The statistics of the LLMs used in this paper. $\dagger$: Non-embedding parameters according to Biderman et al. (2023).
@@
Appendix F Experimental Settings
F.1 Backbones
@@
We use the Pythia suite Biderman et al. (2023) and other popular open-source models, including LLaMa and Vicuna, for our experiments. Pythia is based on GPT-NeoX Black et al. (2022) and includes 8 model sizes and 154 pre-training checkpoints, facilitating research in interpretability and learning dynamics. The statistics of the 9 LLMs used in this paper are summarized in Table 10. We download the pre-trained weights from Huggingface (Wolf et al., 2019).
F.2 Implementation Details
@@
We sort the concepts alphabetically and shuffle the order using random seed 1. The maximum input and output lengths are set to 32 and 10, respectively. The batch size is 32, and the learning rate for the LLMs is $1\times 10^{-5}$. We use the AdamW optimizer Loshchilov and Hutter (2018). For LLMs with more than 1B parameters, we use A800 GPUs, while smaller LLMs run on RTX3090 GPUs. Each experiment is repeated three times, and we report the average and standard deviations. Additionally, we search for the best hyper-parameters for each baseline method.
For the experiment in Section 3.2, we use “gpt-3.5-turbo” and “gpt-4-turbo” for GPT-3.5 and GPT-4, respectively. Given the high cost of evaluating the entire Concept-1K dataset, we randomly sample 500 instances for both GPT-3.5 and GPT-4. The outputs and targets of GPT-3.5 and GPT-4 on these 500 instances are provided in the supplementary material.
F.3 Input Prompt
We use the following input prompt for training and testing Concept-1K:
Question: {Question}
Short Answer: {Answer} ,
where {Question} and {Answer} represent the question and the target output, respectively. For the experiment in Section 3.2, the prompts are provided in Tables 11 and 12.
Table 11: Prompt for In-Context Learning with 1 shot demonstration. {Question i} and {Answer i} represent the question and the target output of the $i$-th demonstration training sample respectively. {Test Question} represents the test question.
I will provide some knowledge as follows:
Question: {Question 1}
Short Answer: {Answer 1}
Please answer the following question according to the above knowledge:
Question: {Test Question}
Short Answer:
Table 12: Prompt for In-Context Learning with 5 shot demonstrations. {Question i} and {Answer i} represent the question and the target output of the $i$-th demonstration training sample respectively. {Test Question} represents the test question.
I will provide some knowledge as follows:
Question: {Question 1}
Short Answer: {Answer 1}
Question: {Question 2}
Short Answer: {Answer 2}
Question: {Question 3}
Short Answer: Answer 3
Question: {Question 4}
Short Answer: {Answer 4}
Question: {Question 5}
short Answer: {Answer 5}
Please answer the following question according to the above knowledge:
Question: {Test Question}
Short Answer:
Appendix G Introduction of Baseline Methods
The introduction of the baseline methods is as follows:
  • SEQ: Sequential fine-tuning (SEQ) is considered the lower bound of incremental learning.
  • REPLAY: Experience replay stores representative old samples and jointly optimizes both old and new samples when learning new tasks. This is a practical and popular technique in incremental learning.
  • EWC Kirkpatrick et al. (2017): Elastic Weight Consolidation (EWC) is a regularization-based method where the weight of each parameter is determined by the diagonal of the Fisher information matrix. The regularization loss weight is searched within {5×$\times 10^{3}$, 1×$\times 10^{4}$, 5×$\times 10^{4}$, 1×$\times 10^{5}$, 1×$\times 10^{6}$, 1×$\times 10^{7}$}.
  • LAMOL Sun et al. (2020): LAMOL trains LLMs with question-answering and generative objectives, generating pseudo-samples before learning each new task for data replay. The generation loss weight is $\lambda=0.25$, and the proportion of pseudo-samples is $\gamma=0.20$. There are two variants: LAMOL_t and LAMOL_g, differing by whether a task-specific token is used for generation.
  • L2KD Chuang et al. (2020): L2KD adds a knowledge distillation target based on LAMOL, with the teacher model trained from scratch. We implemented the word-level variant as it performs best on text classification tasks.
  • LAMOL_KD Zheng et al. (2023b): LAMOL_KD utilizes knowledge distillation based on LAMOL_t. Unlike L2KD, the teacher model in LAMOL_KD is trained on all previous tasks. New data are used to learn the LAMOL objectives, and pseudo data are used for word-level knowledge distillation as a regularization term.
  • PCLL Zhao et al. (2022): PCLL combines the concepts of variational autoencoders and word-level knowledge distillation with the objectives of LAMOL.
  • LFPT5 Qin and Joty (2022): LFPT5 learns only soft prompts for each new task. The training objective is the same as LAMOL. The number of soft prompt tokens is 10.
  • LoRA Hu et al. (2021): LoRA trains a small proportion of parameters of LLMs. We set the rank $r=8$ and the scaling parameter $\alpha=16$. We use the LoRA implementation from the PEFT library Mangrulkar et al. (2022).
Some IL methods are not compared as they are not applicable in the IIL scenario. For example, Progressive Prompt Razdaibiedina et al. (2023) requires task IDs during both training and inference stages. VAG Shao et al. (2023b) requires storing the vocabulary of class labels and does not apply to generation tasks without class labels. Additionally, prompt-based IL methods such as L2P Wang et al. (2022b) are not suitable for generation tasks.
(a) After Task 1
(b) After Task 2
(c) After Task 4
(d) After Task 6
(e) After Task 8
(f) After Task 10
Figure 7: The visualization of the memorized knowledge related to the “Brain-Computer Interface” in IL. The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding training samples are answered correctly.
(a) After Task 1
(b) After Task 2
(c) After Task 4
(d) After Task 6
(e) After Task 8
(f) After Task 10
Figure 8: The visualization of the generalized knowledge related to the “Brain-Computer Interface” in IL. The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding test samples are answered correctly.
Appendix H Visualization of One Concept
We visualize the memorized and generalized knowledge of the concept “Brain-Computer Interface” in Figures 7 and 8. In each graph, the center node represents the concept “Brain-Computer Interface”. The linked and unlinked edges indicate whether the corresponding training or test samples are answered correctly.
For example, in Figure 7a, the edge between “Brain-Computer Interface” and “Signal Processing” signifies that the LLM correctly outputs the target answer “Signal Processing” when the question is the training sample “What subevent occurs in a Brain-Computer Interface?”. In Figure 7b, the edge between “Brain-Computer Interface” and “Signal Processing” is missing, indicating that the LLM fails to provide the correct target answer.
Figure 7 shows that even when LLMs can memorize all the knowledge, they tend to first forget more complex knowledge, such as “(Brain-Computer Interface, UsedFor, Controlling Computers With Thought)”. Conversely, some common knowledge, such as “(Brain-Computer Interface, Requires, Brain Signals)”, “(Brain-Computer Interface, Has Property, Innovative)”, and “(Brain-Computer Interface, Motivated By Goal, Accessibility)”, remains robust and is not forgotten after learning 10 tasks. This indicates that certain knowledge is easier to memorize, generalize, and retain.
Further exploration at the concept-level knowledge in IL is left for future work. We also encourage future studies to utilize the provided Concept-1K dataset for a fine-grained analysis of the memorization and generalization dynamics in IL.
Appendix I Additional Experimental Results